# Detection and Measurement of Syntactic Templates in Generated Text

### v1

@@ Abstract
Recent work on evaluating the diversity of text generated by LLMs has focused on word-level features. Here we offer an analysis of syntactic features to characterize general repetition in models, beyond frequent $n$-grams. Specifically, we define syntactic templates and show that models tend to produce templated text in downstream tasks at a higher rate than what is found in human-reference texts. We find that most (76%) templates in model-generated text can be found in pre-training data (compared to only 35% of human-authored text), and are not overwritten during fine-tuning processes such as RLHF. This connection to the pre-training data allows us to analyze syntactic templates in models where we do not have the pre-training data. We also find that templates as features are able to differentiate between models, tasks, and domains, and are useful for qualitatively evaluating common model constructions. Finally, we demonstrate the use of templates as a useful tool for analyzing style memorization of training data in LLMs.
@@ 1 Introduction
An open question about LLMs is what patterns such models learn from pre-training data (Goldberg, 2019; Petroni et al., 2019; Bender et al., 2021; Chen et al., 2024), and whether these patterns generalize across downstream tasks and datasets (Hupkes et al., 2023). While prior work has focused on the quality of generation (Zhang et al., 2019; Dou et al., 2022; Kryściński et al., 2020), and more recently on text generation novelty (McCoy et al., 2023), there has been limited work on characterizing the sorts of patterns that are learned by LLMs.
Consider, for instance, the text from OLMo-Instruct in Figure 1, which is sampled from a corpus of generated movie review summaries. This was produced by prompting the model to summarize a collection of human-written movie reviews: “The Last Black Man in San Francisco is a poignant, beautifully shot film […] creates a unique and intense viewing experience […]”. In this example, we find a total of 35 repeated sequences of part-of-speech (POS) tags of lengths $n=5$ to $8$. While the generated text was not seen in OLMo’s training data, we find that 33 out of the 35 (95%) sequences appear in the pre-training data. As such, while the generated text itself is novel, it relies on common syntactic sequences learned from the training data.
Figure 1: Sample movie meta-reviews generated by OLMo-7B (top) and Mistral-7B (bottom) by prompting the Rotten Tomatoes dataset. Templates appear at varying rates, and differ across models. We extract templates from the entire corpus of generated text for each model, and match the text to the part-of-speech templates (highlighted), following by the frequency of each template.
In this work we quantify and measure LLMs’ usages of repetitive sequences in text generation. We introduce and focus on syntactic templates, namely POS sequences, a syntactic abstraction over texts that showcase diverse sets of patterns. We begin by establishing the connection between frequent templates in the training data and models’ reliance on such templates during generation. Using this finding, we draw a direct connection between training data template statistics and the tendency of models trained on this data to generate these templates. We then evaluate several closed-source models and the syntactic templates they are prone to generating, which teaches us about properties of their training data.
We start by introducing syntactic templates, and defining methods for detecting and measuring such templates in generated texts (§3). Given this definition, we seek to answer the following questions:
RQ1 How templated are outputs generated by instruction-tuned LLMs?
@@
In all, we evaluate eight models on three different tasks (§5). We show how training data templates are memorized and subsequently generated by models trained on them (§6). We then show how such insights allow one to draw conclusions about the training data used by closed models in a downstream summarization task (§7). Finally, we show that our metrics can also be used as a softer version of memorization. For instance, while Carlini et al. (2022) estimates that 1% of texts to be memorized, we find between 0.8 - 3.1% more soft-memorized texts over verbatim memorization, often by replacing numbers and generating synonyms (§8).
@@ 2 Related Work
Past evaluations of diversity in LLM outputs have primarily focused on token-level diversity (Montahaei et al., 2019; Bache et al., 2013). Diverse sampling strategies have been introduced to address the lower token diversity observed in neural text generation (Holtzman et al., 2020; Roberts et al., 2020), however it is unclear whether such sampling strategies also increase the diversity of the syntactic structure in LLMs. Beyond lexical diversity, Padmakumar and He (2023) extend definitions for measuring content diversity, which has broad applications in downstream tasks such as summarization and creative story generation. Recent work has quantified the drop in generated-text diversity specifically relative to the RLHF training process, however this again focused primarily on token level diversity (Kirk et al., 2024). Our work aligns more closely with the first body of work; we measure the syntactic structure of text rather than its semantic content. Most similar to our work is Bär et al. (2012), which broadly evaluates text repetition metrics at the stylistic, content, and lexical level.
@@ 3 Detecting Syntactic Templates
Our goal is to propose an abstract representation of texts, that is able to capture more subtle repetitions than mere text memorization. Repeated runs of literal tokens may not be sufficient for describing such redundancy nor why a summary produced by, e.g. ChatGPT, might seem familiar.
Focusing on syntactic patterns rather than tokens allows us to capture such repetitions. For example, a pattern consisting of the part-of-speech sequence DT JJ NN IN DT JJ NN will match to phrases in movie reviews (“a romantic comedy about a corporate executive”) and in news summarization (“a humorous insight into the perceived class”), despite these sentences having only one token overlap.
@@ 3.2 Extracting Syntactic Templates from Text
We operationalize the definition in 3.1 as parts-of-speech (POS). For sub-sequences of POS we consider templates of length $n\in\{4,5,6,7,8\}$. Templates are characterized by their high frequency across the texts in a given corpus (e.g., one comprising texts generated by a particular LLM). Practically, we choose $\tau$ relative to the sample size. For Rotten Tomatoes we retain the top 100 most common template where the least frequent template appears 4 times in the dataset.
To extract templates we use diversity (Shaib et al., 2024), [^1] https://pypi.org/project/diversity/ a library providing tools to evaluate token and POS diversity in a dataset. We use this tool to first tag all tokens in a corpus with their corresponding POS tags, then search for the top 100 most frequent $n$-grams across these tags. diversity uses the NLTK POS tagger, [^2] https://www.nltk.org/api/nltk.tag.html which relies on the Penn Treebank set of 36 tags (Taylor et al., 2003). After tagging, we use the tool to return frequent $n$-grams, and the corresponding matched text.
@@ 3.3 Metrics for Measuring Templates
Our goal for extracting templates is to assess and characterize different levels of repetition in LLMs. We use templates to measure three metrics (1) the diversity of the POS tags that are generated, (2) the fraction of texts generated with a template, and (3) the number of templates that appear within each text. We now elaborate on each one.
@@
At the most granular level, we are interested in quantifying the $n$-gram diversity of the POS tag sequences present in the text. Lossless text compression algorithms—such as gZip—are optimized to detect repeated characters in sequences, and rely on this to compress documents without any loss of information. Prior work shows that using gZip to calculate a compression ratio (CR) can provide an efficient measure for capturing lexical diversity, specifically $n$-gram repetition (Shaib et al., 2024).
@@
$\text{CR}(f(T\oplus))=\frac{|f(T\oplus)|}{\text{compressed }|f(T\oplus)|}$ (1)
@@
Template Rate and Count
We measure the fraction of texts that contain at least 1 template to quantify how frequently templates appear across an entire corpus of text.
@@
Where $T$ is a concatenated sequence of tokens forming a corpus for a particular text source, and $t$ the individual text.
@@ 4 Experimental Setup
We first evaluate the incidence of templated text in two open-ended generation tasks using OLMo-7B Instruct (Groeneveld et al., 2024), a completely open source model that released model training checkpoints and its training data. This allows us to evaluate templates in its training datasets: Dolma (Soldaini et al., 2024), Tulu-V2 (Ivison et al., 2023), and Ultra-feedback (Cui et al., 2023).
@@
Decoding
@@
4.1 Fine-tuned (Instruction) Models
We experiment with a total of 8 instruction-tuned models. We use Mistral (Instruct, 7B; Jiang et al. 2023), Alpaca (7B, 13B; Taori et al. 2023; Wang et al. 2022; Touvron et al. 2023a). In addition, 5 models are further trained on human preferences: OLMo (Instruct, 7B; Groeneveld et al. 2024), Llama-2 (Chat-HF, 7B-70B; Touvron et al. 2023b), and Llama-3 (Instruct, 70B AI@Meta 2024).
@@ 4.2 Tasks and Datasets
To evaluate intrinsic template behaviour we evaluate over an open-ended generation task. In the first setting, we sample generations from the model given only a special token denoting beginning of sequence ([BOS]). In the second, we randomly sample 100 tokens from Dolma and use these tokens to prompt further open-ended generation from the model.
@@
LLMs are increasingly used to create synthetic training datasets (e.g., Wang et al. 2022). We evaluate templates in Cosmopedia, a synthetic dataset generated by prompting Mixtral-8x7B-Instruct with instructions to produce text relating to textbooks, blogposts, stories, posts and WikiHow articles (Ben Allal et al., 2024). We prompt OLMo-7B with the Cosmopedia instructions and evaluate the resulting generations.
@@
is a common benchmark for long text generation. We evaluate models on a handful of summarization datasets, including single- and multi-document tasks. Such datasets allow us to study templates in longer sequences that would not be evident in tasks where only a few tokens are generated. For general English-language tasks, we generate summaries and reviews over news (CNN/Daily Mail; Nallapati et al. 2016), movies (Rotten Tomatoes; Leone 2020), and books (BooookScore; Chang et al. 2023).
@@ 5 Templates in Model-Generated Text
We first evaluate OLMo-7B Instruct on 4 tasks: open-ended generation, synthetic data generation, and summarization, using both greedy and varying temperature and top-$p$ sampling strategies (RQ 1).
Table 1 shows the effect of varying sampling hyperparameters temperature and top-$p$ on the overall diversity of the generated text with OLMo-7B with open-generation and summarization. Varying sampling strategies in the open-generation task results in a higher variance of template rates (74.4% $\pm$ 2.1) compared to templates rates in the summarization task (96.8% $\pm$ 0.6). These results indicate that templatic text in summarization appears in spite of sampling strategies intended to increase (lexical) diversity.
Overall, the rate of templates is much higher in the Rotten Tomatoes dataset than for open-generation, indicating downstream tasks such as summarization, which often entail prompting with instructions, may yield more repetitive structures. [^3] We show the incidence of templates given different instructions in Appendix D and find that, while the rates somewhat vary, there is still a high incidence of templates across all prompts.
Table  2 shows the rate of templates on two additional tasks: Synthetic data generation and data generation with Dolma Cosmopedia results in a higher incidence of templates (99.1%) and templates per token (0.014), compared to Dolma (82.6%, 0.012).
@@ 6.1 Emergence of Templates in Training
We first aim to understand when during training models start to produce templates. We measure this by looking at the perplexity of a set of identified templates across OLMo model checkpoints. Higher perplexity values indicate the templates are assigned low likelihood at that checkpoint. For each model checkpoint, we average the perplexities of templates of length $n=6$ and compare to the perplexities of randomly sampled 6-grams. We calculate the average perplexity for the dataset using:
@@ 6.2 Templates in Pre-training Data
The above finding indicates that templates are seen fairly early on in pre-training. We next measure the incidence and types of templates in the (pre-)training data, and whether they correspond to the templates that models produce.
@@ 7 Templates in Closed-Source Models
With OLMo, we find that most templates are in the pre-training data. We evaluate the incidence of templates in other closed-source models, which we define as models that do not release their training data. Addressing RQ1, we characterize the rates of templatic texts in these models, and posit that templates may be indicators of the pre-training data sources models are trained on.
@@
We report the rate of templates in Table 3, and measure the percentage of outputs that contain at least one template of length $n$. We find that, on average in the Rotten Tomatoes dataset, 95% of outputs contain templates of length $n=6$ across different model types and sizes. This is in contrast to human-written reference and input documents, which contain templates of the same size on average in 38% of cases. We find a similar trend for templates of length $n=[4,8]$ (Figure 5).
@@
The CNN/DM datasets show similar trends, but with lower rates of templates compared to the Rotten Tomatoes dataset. In contrast, the percentage of templates is high for model-generated and human-written references in the Cochrane dataset. This owes to the nature of meta-analysis texts, which are formulaic (Higgins and Green, 2010). [^4] Appendix Table C provides examples of the human-written references
Figure 5 illustrates the rate of templates for each model as the template length grows from length $n=4$ to $8$ for Rotten Tomatoes and Cochrane. For Rotten Tomatoes (and CNN; Appendix B), all models produce templates at higher rates than human-written summaries across all template lengths. With Cochrane, template lengths $\geq 6$ show the majority of models produce higher rates of templates than human authored references.
@@
Model Size

### v2

@@ Abstract
The diversity of text can be measured beyond word-level features, however existing diversity evaluation focuses primarily on word-level features. Here we propose a method for evaluating diversity over syntactic features to characterize general repetition in models, beyond frequent $n$-grams. Specifically, we define syntactic templates (e.g., strings comprising parts-of-speech) and show that models tend to produce templated text in downstream tasks at a higher rate than what is found in human-reference texts. We find that most (76%) templates in model-generated text can be found in pre-training data (compared to only 35% of human-authored text), and are not overwritten during fine-tuning or alignment processes such as RLHF. The connection between templates in generated text and the pre-training data allows us to analyze syntactic templates in models where we do not have the pre-training data. We also find that templates as features are able to differentiate between models, tasks, and domains, and are useful for qualitatively evaluating common model constructions. Finally, we demonstrate the use of templates as a useful tool for analyzing style memorization of training data in LLMs  [^1] https://cshaib.github.io/syntactic_templates/.
@@ 1 Introduction
An open question about large language models (LLMs) is what patterns such models learn from pre-training data (Goldberg, 2019; Petroni et al., 2019; Bender et al., 2021; Chen et al., 2024), and whether the same patterns appear generally across downstream tasks and datasets (Hupkes et al., 2023). While prior work has focused on the quality of generation (Zhang et al., 2019; Dou et al., 2022; Kryściński et al., 2020), and more recently on text generation novelty (McCoy et al., 2023; Merrill et al., 2024), there has been limited work on characterizing the sorts of lexical patterns that are learned by LLMs.
Consider, for instance, the generated text from OLMo-Instruct in Figure 1, which is sampled from a corpus of movie review summaries. This was produced by prompting the model to summarize a collection of human-written movie reviews: “The Last Black Man in San Francisco is a poignant, beautifully shot film […] creates a unique and intense viewing experience […]”. While this generated text was not seen in Dolma Soldaini et al. (2024), OLMo’s pre-training data, we find a total of 35 unique repeated sequences of part-of-speech (POS) tags of lengths $n=5$ to $8$ in the summarized movie reviews. Further, we find that 33 out of the 35 (95%) sequences appear in the pre-training data. As such, while the generated text itself is novel, it relies on common syntactic sequences seen in the training data.
Figure 1: Sample movie meta-reviews generated by OLMo-7B (top) and Mistral-7B (bottom) by prompting the Rotten Tomatoes dataset. Templates appear at varying rates (frequency shown out of 500 generations), and differ across models. We extract templates from the entire corpus of generated text for each model, and match the text to the part-of-speech templates (highlighted), following by the frequency of each template.
In this work, we quantify and measure LLMs’ usages of repetitive sequences in text generation. We introduce and focus on syntactic templates, namely POS sequences, a syntactic abstraction over texts. We seek to answer the following questions:
RQ1 To what extent do outputs generated by instruction-tuned LLMs contain templates?
@@
We start by introducing syntactic templates, and defining methods for detecting and measuring such templates in generated texts (§3). We evaluate eight models on three different tasks (§5). We show how training data templates are memorized and subsequently generated by models trained on them (§6). We then show how such insights allow one to draw conclusions about the training data used by closed models in a downstream summarization task (§7). Finally, we show that our metrics can also be used as a softer version of memorization. For instance, while Carlini et al. (2022) estimates that 1% of texts to be memorized, we find between 0.8 - 3.1% more soft-memorized texts over verbatim memorization, often by replacing numbers and synonyms (§8).
@@ 2 Related Work
Past evaluations of diversity in LLM outputs have primarily focused on token-level diversity (Montahaei et al., 2019; Bache et al., 2013). Diverse sampling strategies have been introduced to address the lower token diversity observed in neural text generation (Holtzman et al., 2020; Roberts et al., 2020), however it is unclear whether such sampling strategies also increase the diversity of the syntactic structure in LLMs. Beyond lexical diversity, Padmakumar and He (2023) extend definitions for measuring content diversity, which has broad applications in downstream tasks such as summarization and creative story generation. Recent work has quantified the drop in generated-text diversity specifically relative to the RLHF training process, however this again focused primarily on token level diversity (Kirk et al., 2024). Our work aligns more closely with the first body of work; we measure the syntactic structure of text rather than its semantic content. Most similar to our work is Bär et al. (2012), which broadly evaluates text repetition metrics at the stylistic, content, and lexical level. Our methods do not address repetition in content but rather focus on extending the characterization of lexical and stylistic repetition with text abstractions in LLMs.
@@ 3 Detecting Syntactic Templates
Our goal is to search for abstract representations of texts to capture more subtle repetitions than mere text memorization. Repeated strings of literal tokens may not be sufficient for describing such redundancy nor why a summary produced by, e.g., ChatGPT, might seem familiar.
Focusing on syntactic patterns rather than tokens allows us to capture such repetitions For example, a pattern consisting of the part-of-speech sequence DT JJ NN IN DT JJ NN will match to phrases in movie reviews (“a romantic comedy about a corporate executive”) and in news summarization (“a humorous insight into the perceived class”), even though these sentences have only one token overlap.
@@ 3.2 Extracting Syntactic Templates from Text
We operationalize the definition in 3.1 as parts-of-speech (POS). For sub-sequences of POS we consider templates of length $n\in\{4,5,6,7,8\}$. Templates are characterized by their high frequency across the texts in a given corpus (e.g., one comprising texts generated by a particular LLM) Practically, we choose $\tau$ relative to the sample size. For Rotten Tomatoes we retain the top 100 most common template where the least frequent template appears 4 times in the dataset.
To extract templates we use diversity (Shaib et al., 2024), [^2] https://pypi.org/project/diversity/ a library providing tools to evaluate token and POS diversity in a dataset. We use this tool to first tag all tokens in a corpus with their corresponding POS tags, then search for the top 100 most frequent $n$-grams across these tags. diversity uses the SpaCy POS tagger, [^3] https://spacy.io/api/tagger which relies on the Penn Treebank set of 36 tags (Taylor et al., 2003). After tagging, we return frequent $n$-grams of POS, the corresponding matched text. Figure 2 illustrates the output of running the template extraction process.
Figure 2: Example templates (left) and matched text (right) returned from the diversity package.
@@ 3.3 Metrics for Measuring Templates
Our goal for extracting templates is to assess and characterize different levels of repetition in LLMs. We calculate three metrics using templates, (1) the diversity of the POS tags that are generated using CR-POS (2) the fraction of texts generated with a template using template rate and count, and (3) the number of templates that appear within each text using templates-per-token. We now elaborate on each one.
@@
At the most granular level, we are interested in quantifying the $n$-gram diversity of the POS tag sequences present in the text. Lossless text compression algorithms—such as gZip—are optimized to detect repeated characters in sequences, and rely on this to compress documents without any loss of information. If a document contains frequent repeated strings, the document will be more compressible, resulting in a larger difference in compressed size relative to the original document size. Shaib et al. (2024) show that using gZip to calculate a compression ratio (CR) can provide an efficient measure for capturing lexical diversity, specifically $n$-gram repetition.
@@
$\text{CR}(f(T\oplus))=\frac{|f(T\oplus)|}{\text{compressed }(|f(T\oplus)|)}$ (1)
@@
Template Rate
We measure the fraction of texts in a corpus (sequence) that contain at least 1 template to quantify how frequently templates appear across an entire corpus.
$TR=\frac{1}{T}\sum_{i=1}^{T}I_{i}$ (2)
Where, $T$ is the sequence of text (corpus), and
$I_{i}$ =$=\begin{cases}1&\text{if text }i\text{ contains at least 1 template}\\ 0&\text{otherwise}\end{cases}$
@@
Where $T$ is a concatenated sequence of tokens forming a corpus for a particular text source, and $t$ the string.
@@ 4 Experimental Setup
4.1 Models
Open Models
We first evaluate the incidence of templated text in two open-ended generation tasks using OLMo-7B Instruct (Groeneveld et al., 2024), a fully open source model that released model training checkpoints and its training data. This allows us to evaluate templates in its training datasets: Dolma (Soldaini et al., 2024), Tulu-V2 (Ivison et al., 2023), and Ultra-feedback (Cui et al., 2023).
@@ 4.1 Models
Fine-tuned (Instruction) Models
We experiment with a total of 8 instruction-tuned models. We use Mistral (Instruct, 7B; Jiang et al. 2023), Alpaca (7B, 13B; Taori et al. 2023; Wang et al. 2022; Touvron et al. 2023a). In addition, 5 models are further trained on human preferences: OLMo (Instruct, 7B; Groeneveld et al. 2024), Llama-2 (Chat-HF, 7B-70B; Touvron et al. 2023b), and Llama-3 (Instruct, 70B Dubey et al. 2024).
4.2 Decoding
@@ 4.3 Tasks and Datasets
To evaluate intrinsic template behaviour we evaluate open-ended generation tasks in two settings. In the first setting, we sample generations from the model given only a special token denoting beginning of sequence ([BOS]). In the second, we randomly sample 100 tokens from Dolma and use these tokens to prompt further open-ended generation from the model.
@@
LLMs are increasingly used to create synthetic training datasets, which are often used to train downstream models (e.g., Wang et al. 2022). We evaluate templates in Cosmopedia, a synthetic dataset generated by prompting Mixtral-8x7B-Instruct with instructions to produce text relating to textbooks, blogposts, stories, posts and WikiHow articles (Ben Allal et al., 2024). We prompt OLMo-7B with the Cosmopedia instructions and evaluate the resulting generations.
@@
Table 1: Compression ratio with POS (CR-POS), average text length and percentage of generated outputs with at least 1 template of size $n=6$, when varying temperature and top_p for OLMo-7B decoding. Arrows indicate higher template rates.
@@
Summarization is a common benchmark for long text generation. We evaluate models on a handful of summarization datasets, including single- and multi-document tasks. Such datasets allow us to study templates in longer sequences that would not be evident in tasks where only a few tokens are generated. For general English-language tasks, we generate summaries and reviews over news (CNN/Daily Mail; Nallapati et al. 2016), movies (Rotten Tomatoes; Leone 2020), and books (BooookScore; Chang et al. 2023).
@@ 5 Templates in Model-Generated Text
We first evaluate OLMo-7B Instruct on 3 tasks: open-ended generation, synthetic data generation, and summarization, using both greedy and varying temperature and top-$p$ sampling strategies (RQ1).
Table 1 shows the effect of varying sampling hyperparameters temperature and top-$p$ on the overall diversity of the generated text with OLMo-7B with open-generation and summarization.
Varying sampling strategies in the open-generation task results in a higher variance of template rates (74.4% $\pm$ 2.1) compared to templates rates in the summarization task (96.8% $\pm$ 0.6). These results indicate that templatic text in summarization appears in spite of sampling strategies intended to increase (lexical) diversity. Overall, the rate of templates is much higher in the Rotten Tomatoes dataset than for open-generation, indicating downstream tasks such as summarization, which often entail prompting with instructions, may yield more repetitive structures. [^4] Note that we show the incidence of templates given different instructions in Appendix D
Table  2 shows the rate of templates on two additional tasks: Synthetic data generation and data generation with Dolma. Cosmopedia results in a higher incidence of templates (99.1%) and templates per token (0.014), compared to Dolma (82.6%, 0.012).
@@ 6.1 Emergence of Templates in Training
We first aim to understand when during training models start to generate templates. We measure the perplexity of matched texts from a set of previously extracted templates (following §3.2) across OLMo’s checkpoints. Higher perplexity values indicate the templates are assigned low likelihood at that checkpoint.
For each model checkpoint, we average the perplexities of templates of length $n=6$ and compare to the perplexities of randomly sampled 6-grams. We calculate the average perplexity for the dataset using:
@@ 6.2 Templates in Pre-training Data
The lower perplexities in the above finding indicates that templates are seen fairly early on in pre-training compared to non-templated sequences of training data text. We next measure the incidence and types of templates in the pre-training data, and whether they correspond to the templates that models produce.
@@ 7 Templates in Closed-Source Models
With OLMo, we find that 75% of templates are found at high frequencies in the pre-training data (§6.2). Most available models however do not release their pre-training data. Here we evaluate the incidence of templates in other closed-source models, which we define as models that do not release their training data. Addressing RQ1, we characterize the rates of templatic texts in these models, and posit that templates may be indicators of the pre-training data sources models are trained on.
@@
We report the template rate and templates-per-token that appear in text generated by models in three summarization tasks: movie reviews, biomedical evidence, and news (Table 3).
We find that, on average in the Rotten Tomatoes dataset, 95% of outputs contain templates of length $n=6$ across different model types and sizes. This is in contrast to human-written reference and input documents, which contain templates of the same size on average in 38% of cases. We find a similar trend for templates of length $n=[4,8]$ (Figure 6).
@@
The CNN/DM datasets show similar trends, but with lower rates of templates (average 89.6% contain templates) compared to the Rotten Tomatoes dataset. In contrast, the percentage of templates is high for model-generated (average 88.3%) and human-written references (83.3%) in the Cochrane dataset. This owes to the nature of meta-analysis texts, which are formulaic (Higgins and Green, 2010). [^5] Table C in the appendix provides examples of the human-written references
Figure 6 illustrates the rate of templates for each model as the template length grows from length $n=4$ to $8$ for Rotten Tomatoes and Cochrane. For Rotten Tomatoes (and CNN; Appendix B), all models produce templates at higher rates than human-written summaries across all template lengths. With Cochrane, template lengths $\geq 6$ show the majority of models produce higher rates of templates than human authored references. This indicates that differences between templatedness in human-authored references and LLM summaries surface only at longer template lengths.
@@
Effect of Model Size on Template Rates